# ResearchTown: Simulator of human research community

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable potential in scientific domains, yet a fundamental question remains unanswered: Can we simulate human research communities with LLMs? Addressing this question can deepen our understanding of the processes behind idea brainstorming and inspire the automatic discovery of novel scientific insights. In this work, we propose ResearchTown, a multi-agent framework for research community simulation. Within this framework, the human research community is simplified and modeled as an agent-data graph, where researchers and papers are represented as agent-type and data-type nodes, respectively, and connected based on their collaboration relationships. We also introduce TextGNN, a text-based inference framework that models various research activities (e.g., paper reading, paper writing, and review writing) as special forms of a unified message-passing process on the agent-data graph. To evaluate the quality of the research simulation, we present ResearchBench, a benchmark that uses a node-masking prediction task for scalable and objective assessment based on similarity. Our experiments reveal three key findings: (1) ResearchTown can provide a realistic simulation of collaborative research activities, including paper writing and review writing; (2) ResearchTown can maintain robust simulation with multiple researchers and diverse papers; (3) ResearchTown can generate interdisciplinary research ideas that potentially inspire novel research directions.
@@ 1 Introduction
LLMs have proved themselves to be powerful copilots in scientific research (AI4Science & Quantum, 2023), demonstrating their great potential for accelerating and automating scientific discovery. Despite the promising finding, there remains a more ambitious question to answer: Can we simulate human research community with LLMs? Answering such a question has multiple benefits: (1) simulating research communities helps understand the underlying process behind the discovery of existing research ideas; (2) it can further help democratize and accelerate the idea discovery process.
@@
However, simulating the human research community is challenging, as it involves leveraging multiple LLM agents to interact with heterogeneous and research-related data. While existing multi-agent LLM frameworks have been successfully applied to areas like social simulation (Zhou et al., 2023; Gao et al., 2023) and game simulation (Hua et al., 2023; Xu et al., 2023), they are not well-suited for simulating research communities due to the complexity of collaborative research activities like paper writing and review writing. Although recent efforts have explored research automation using LLMs, these frameworks are typically limited to specific research tasks, such as idea generation (Girotra et al., 2023) or code experimentation (Huang et al., 2024), or focus on simulating the workflow of a single researcher (Lu et al., 2024). These frameworks cannot simulate collaborative research activities where researchers from diverse backgrounds and expertise work together to brainstorm ideas, have discussions, and review papers—processes that are fundamental to modern human research.
Research community as graph. Our key observation is that the deeply interconnected research community can be naturally represented as graphs. Indeed, similar graph structures like citation networks (Newman, 2001) and academic social networks (Tang et al., 2008) have been extensively studied within data mining research, with proven values in applications such as citation prediction (Holm et al., 2020), recommendation (West et al., 2016) and community detection (Yang & Leskovec, 2012). Introducing LLMs to a graph-structured research community can extend these classic works from static analysis to dynamic simulation and real-time forecasting.
Novel framework for research simulation. In this work, we propose ResearchTown, a simulator of the human research community. To bridge the gap between existing multi-agent simulation frameworks and the complexity of research activities, we propose a graph-based framework, inspired by the message-passing mechanism in Graph Neural Networks (GNNs), for multi-agent simulation. Concretely, as shown in Figure 1, we propose a new concept of agent-data graph with 2 generic types of nodes: agent nodes, suitable for entities like agents, and data nodes, suitable for entities such as papers, reviews, and blogs. Agent-data graphs are unique from standard heterogeneous graphs; here, the key conceptual difference between agent and data nodes is that an agent node can be considered a function over data nodes. To inference on agent-data graphs, we propose a TextGNN framework where message-passing processes are defined based on text-form information processing with LLMs, thanks to their strong in-context learning (Wei et al., 2023) and reasoning (Lee et al., 2024) ability. We apply the proposed agent-data graph and TextGNN to the research simulation. Here, a research community can be regarded as a special form of agent-data graph, called community graph, with research agents and research papers as two types of nodes, and we consider three types of edges (review, author, and cite) in the graph. Different community activities, such as paper writing and review writing, can be modeled as special message-passing processes on the community graph.
@@
Figure 1: Abstracting and simplifying human research community as an agent-data graph, i.e. community graph. An agent-data graph has researchers as agent nodes and blogs, codebases, and papers as data nodes. Without loss of generality, we abstract it into a simplified version with only researcher and paper nodes and focus on critical research tasks including paper reading, paper writing, and review writing. Each data node has a hidden state $h_{u}$ and each agent node is paired with an agent function $f_{v}(\cdot)$ and a hidden state $h_{v}$.
@@
Novel evaluation for research simulation. With ResearchTown for research simulation, a further research question is to evaluate the quality of that. Prior works primarily use human evaluation with breakdown metrics like novelty, excitement, feasibility, and expected effectiveness (Si et al., 2024; Hu et al., 2024b). These approaches inevitably suffer from subjectiveness and high costs. In our work, graph-based ResearchTown naturally provides a similarity-based evaluation method, by masking a given paper node in the community graph and evaluating if a simulator can reconstruct the masked nodes. Such a definition gets rid of high-quality human annotations, making it scalable and objective. With the help of such node masking prediction task, we build a benchmark called ResearchBench with 1,000 paper writing tasks and 200 review writing tasks for systematic evaluation.
@@
Main discoveries. Based on the evaluation results from ResearchBench, we highlight three key findings: (1) ResearchTown accurately simulates collaborative research activities, achieving a similarity score of 0.67 for paper writing and 0.49 for review writing, as measured by state-of-the-art text embedding models; (2) ResearchTown demonstrates robustness in research simulation, showing continuous improvement when more researchers or reviewers are added and maintaining consistent performance when including unrelated papers; (3) ResearchTown inspires interdisciplinary research, generating innovative ideas that combine insights from NLP, criminology, and astronomy and does not exist in the real-world research.
@@
Stressing ethical concerns. As our work targets conducting automatic research and simulating the human research community, multiple ethical concerns including potential research plagiarism appear. These ethical concerns are addressed in detail in Appendix §B.
@@ 2 Additional related work
Graphs with text attributes. In real-world graph tasks, nodes often have textual attributes to carry richer information, forming text-attributed graphs (TAGs) (Yang et al., 2021; He et al., 2023). Previous work on TAGs mainly treats LLMs as tools for understanding text attributes and aims at achieving co-training LLMs and GNNs (Zhao et al., 2023; Chen et al., 2024). In contrast, our approach incorporates agent nodes into the graph, enabling text-based message passing between agent nodes and data nodes. Furthermore, while previous TAG research mainly focuses on node prediction and link prediction tasks (Yan et al., 2023), ResearchTown extends this scope by targeting both the reconstruction of existing nodes and the prediction of new, non-existent nodes.
@@ 3 Agent-data graph for multi-agent LLMs
Definition of agent-data graphs. To initiate our discussion, we formally define the proposed agent-data graph. An agent-data graph is a special type of heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\mathcal{V}_{a}\cup\mathcal{V}_{d}$ is the node set consisting of two types of nodes, agent nodes and data nodes, and $\mathcal{E}=\mathcal{E}_{aa}\cup\mathcal{E}_{ad}\cup\mathcal{E}_{dd}$ is the edge set consisting of three types of relations, agent-agent, data-data, and agent-data interactions. Here, each data node $v\in\mathcal{V}_{d}$ comes with attributes, e.g., a piece of text, $\mathbf{x}_{v}$; each agent node $u$ is accompanied with an agent function, e.g., an LLM $f_{u}(\cdot)$ with its prompt template and the profile. Each agent function is responsible for two types of tasks: message generation and message aggregation. More details about agent functions are in Appendix §C.3. Without loss of generality, we assume the data nodes have text attributes, and leave the extension of our work to multi-modal information, e.g., images, audio, and videos, to future works.
@@
Example of agent-data graphs. Figure 1 shows a concrete example of the agent-data graph. Its definition could be extended to more node types (e.g., codebase, blogs) and edge types (e.g., attend, post, commit). Typically, the appearance of one blog post can be directly related to multiple researchers, papers, and other blog posts. Therefore, such entities are directly connected with it.
@@ 4 Building TextGNN on agent-data graphs
$\mathbf{m}_{u}^{(k)}=\textsc{Msg}^{(k)}(\mathbf{h}_{u}^{(k-1)})\quad\quad\mathbf{h}_{v}^{(k)}=\textsc{Agg}^{(k)}\big{(}\mathbf{h}_{v}^{(k-1)},\{\mathbf{m}_{u}^{(k)}\mid u\in\mathcal{N}(v)\}\big{)}$ (1)
@@
Message passing in TextGNN. Following the message-passing process in the standard GNN, we now define a general form of the aggregation function to describe the text-based message-passing process on an agent-data graph $\mathcal{G}$. The key difference between a standard GNN and a TextGNN is that all the hidden states in standard GNN are defined in the embedding space ($\mathbf{h}_{v}\in\mathbb{R}^{d}$) while those in TextGNN are defined in the text space ($\mathbf{h}_{v}\in\Sigma^{*}$). In a TextGNN, we first set the initial hidden states for data nodes $\mathbf{h}_{v}^{(0)}=\mathbf{x}_{v}$, where $\mathbf{x}_{v}$ are text attributes, and the initial hidden states for agent nodes is empty $\mathbf{h}_{u}^{(0)}=\emptyset$. Next, we design a general form of message passing function that handles three distinctive types of interactions, agent-agent $\mathcal{E}_{aa}$, agent-data $\mathcal{E}_{ad}$, and data-data $\mathcal{E}_{dd}$.
Specifically, the $k$-th TextGNN layer for an agent node $u\in\mathcal{V}_{a}$ can be written as
$\begin{split}\mathbf{h}_{u}^{(k)}&=\textsc{Agg}^{(k)}\big{(}f_{u}(\cdot),\mathbf{h}_{u}^{(k-1)},\{f_{a}(\cdot),\mathbf{h}_{a}^{(k-1)}\mid(u,a)\in\mathcal{E}_{aa}\},\{\mathbf{h}_{d}^{(k-1)}\mid(u,d)\in\mathcal{E}_{ad}\}\big{)}\\ &=f_{u}\left(\left[\mathbf{h}_{u}^{(k-1)},\left\{f_{a}\big{(}\big{[}\mathbf{h}_{a}^{(k-1)},\mathbf{h}_{u}^{(k-1)},\mathbf{h}_{d}^{(k-1)}\big{]}\big{)}\mid(u,a)\in\mathcal{E}_{aa},(u,d)\in\mathcal{E}_{ad}\right\}\right]\right)\end{split}$ (3)
where []delimited-[][$\cdot$][ ⋅ ] is the concatenation function between texts to fill in the prompt template, $\mathbf{h}_{v}^{(k)}$ represents the hidden states of the $k$-th layer of $v\in\mathcal{V}$, $f_{a}(\cdot)$ represents the agent function paired with the agent node in the neighborhood and $f_{u}(\cdot)$ represents the agent function paired with the target agent node.
Similarly, the $k$-th TextGNN layer for a data node $v\in\mathcal{V}_{d}$ can be written as
$\begin{split}\mathbf{h}_{v}^{(k)}&=\textsc{Agg}^{(k)}\big{(}\mathbf{h}_{v}^{(k-1)},\{f_{a}(\cdot),\mathbf{h}_{a}^{(k-1)}\mid(v,a)\in\mathcal{E}_{ad}\},\{\mathbf{h}_{d}^{(k-1)}\mid(v,d)\in\mathcal{E}_{dd}\}\big{)}\\ &=f_{g}\left(\left[\mathbf{h}_{v}^{(k-1)},\left\{f_{a}\big{(}\big{[}\mathbf{h}_{a}^{(k-1)},\mathbf{h}_{v}^{(k-1)},\mathbf{h}_{d}^{(k-1)}\big{]}\big{)}\mid(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\right\}\right]\right)\end{split}$ (4)
@@ 5 ResearchTown: Applying TextGNN to community graph
Overview of ResearchTown. Based on the definition of the TextGNN and the agent-data graph, we can apply them to research community simulation to describe different research activities, where each type of activity can be regarded as a different instantiation of the TextGNN layer. The overall ResearchTown simulation process takes a set of papers as input and takes a generated paper and a generated review corresponding to that paper as outputs. We first describe the concept of a community graph as the instantiation of an agent-data graph in research simulation. Then, we describe the specific TextGNN layers that are used to model each type of research activity.
@@
Agent-data graph for research community modeling - community graph. We adapt agent-data graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to research community simulation, which we named as community graph. As is shown in Figure 2, here, the agent nodes $\mathcal{V}_{a}$ are researchers, and the data nodes $\mathcal{V}_{d}$ are research papers. We consider edge set $\mathcal{E}_{dd}$ as paper citations, edge set $\mathcal{E}_{ad}$ as a researcher authors a paper and/or a researcher has the expertise to review it. We omit the edge set $\mathcal{E}_{aa}$ to simplify the framework since oftentimes author collaboration relations can be captured by 2-hop $\mathcal{E}_{ad}$ authorship relations.
Figure 2: ResearchTown simulation as TextGNN inference on the community graph. We consider the process of adding a non-existent paper node into the community graph including three stages: paper reading to insert agent nodes, paper writing to insert data nodes, and review writing for deciding whether to drop generated nodes or not.
@@
TextGNN for research activity simulation. Based on the constructed community graph, we further identify the key types of research activities where TextGNN can be used for simulation. Specifically, we split the research simulation process includes three critical stages: (1) paper reading (2) paper writing (3) review writing. We believe these stages are crucial in the research community and each stage relies on the output of the previous stage as the input. We provide a detailed description for each stage and the corresponding TextGNN layer definition below.
@@
Stage 1: Paper reading. Reading papers to collect insights is a necessary process for initializing a research project. In the community graph, the paper reading process can be described as inserting a new agent node to the community graph and aggregating its neighborhood information based on Equation 3. Here, the new agent profile is non-existent before reading a collection of papers, and the profile is created after the paper reading process, making the TextGNN layer unique. Concretely, by adapting Equation 3, the TextGNN layer for paper reading can be written as:
@@
$\begin{split}\mathbf{h}_{u}&=\textsc{Agg}\big{(}f_{u}(\cdot),\{\mathbf{h}_{d}\mid(u,d)\in\mathcal{E}_{ad}\}\big{)}\\ &=f_{u}\left(\left[\left\{\mathbf{h}_{d}\mid(u,d)\in\mathcal{E}_{ad}\right\}\right]\right)\end{split}$ (5)

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable potential in scientific domains, yet a fundamental question remains unanswered: Can we simulate human research communities with LLMs? Addressing this question can deepen our understanding of the processes behind idea brainstorming and inspire the automatic discovery of novel scientific insights. In this work, we propose ResearchTown, a multi-agent framework for research community simulation. Within this framework, the human research community is simplified as an agent-data graph, where researchers and papers are represented as agent-type and data-type nodes, respectively, and connected based on their collaboration relationships. We also introduce TextGNN, a text-based inference framework that models various research activities (e.g., paper reading, paper writing, and review writing) as special forms of a unified message-passing process on the agent-data graph. To evaluate the quality of the research community simulation, we present ResearchBench, a benchmark that uses a node-masking prediction task for scalable and objective assessment based on similarity. Our experiments reveal three key findings: (1) ResearchTown can provide a realistic simulation of collaborative research activities, including paper writing and review writing; (2) ResearchTown can maintain robust simulation with multiple researchers and diverse papers; (3) ResearchTown can generate interdisciplinary research ideas that potentially inspire pioneering research directions.
Machine Learning, ICML
[…]
footnote
@@ 1 Introduction
LLMs have proved to be powerful copilots in scientific research (AI4Science & Quantum, 2023), demonstrating their great potential for accelerating scientific discovery. Despite the promising finding, a more ambitious question remains: Can we simulate the human research community with LLMs? Answering such a question has multiple benefits: (1) simulating the human research community helps understand the underlying process behind the discovery of existing research ideas; (2) it can further help democratize and accelerate the discovery process of new research ideas.
@@
However, simulating the human research community is challenging, as it involves leveraging multiple LLM agents to interact with complex research data. While existing multi-agent LLM frameworks have been successfully applied to areas like social simulation (Zhou et al., 2023; Gao et al., 2023) and game simulation (Hua et al., 2023; Xu et al., 2023), they are not well-suited for simulating research communities due to the complexity of collaborative research activities like paper writing and review writing. Although recent efforts have explored research automation using LLMs, these frameworks are typically limited to specific research tasks, such as idea generation (Girotra et al., 2023; Baek et al., 2024) or code experimentation (Huang et al., 2024), or focus on simulating single-agent workflows (Lu et al., 2024). These frameworks cannot simulate collaborative research activities where researchers with diverse backgrounds work together to brainstorm ideas, review papers, etc—processes that are fundamental to modern human research.
@@
Figure 1: Abstracting and simplifying human research community as an agent-data graph, i.e., community graph. An agent-data graph has researchers as agent nodes and blogs, codebases, and papers as data nodes. Without losing generality, we abstract it into a simplified version with only researcher and paper nodes and focus on critical research tasks, including paper reading, paper writing, and review writing. Each data node has a hidden state $h_{u}$, and each agent node is paired with an agent function $f_{v}(\cdot)$ and a hidden state $h_{v}$.
Research community as graph. Our key observation is that the deeply interconnected research community can be naturally represented as graphs. Indeed, similar graph structures like citation networks (Newman, 2001) and academic social networks (Tang et al., 2008) have been extensively studied within data mining research, with proven values in applications such as citation prediction (Holm et al., 2020), recommendation (West et al., 2016), and community detection (Yang & Leskovec, 2012). However, introducing LLMs to a graph-structured research community can extend these previous works from prediction and analysis with existing data to dynamic simulation and real-time forecasting.
Novel framework for research simulation. In this work, we propose ResearchTown, a simulator of the human research community. To bridge the gap between existing multi-agent simulation frameworks and the complexity of research activities, we propose a graph-based framework, inspired by the message-passing mechanism in Graph Neural Networks (GNNs), for multi-agent simulation. Concretely, as shown in Figure 1, we propose a new concept of agent-data graph with 2 generic types of nodes: (1) agent nodes, suitable for entities like agents; (2) data nodes, suitable for entities such as papers, reviews, and blogs. Agent-data graphs are unique from standard heterogeneous graphs; here, the key conceptual difference between agent and data nodes is that an agent node can be considered a function over data nodes. To inference on agent-data graphs, we propose a TextGNN framework where message-passing processes are defined based on text-form information processing with LLMs, thanks to their strong in-context learning (Wei et al., 2023) and reasoning (Lee et al., 2024) ability. We apply the proposed agent-data graph and TextGNN to the research simulation. Here, a research community can be regarded as a special form of agent-data graph, called community graph, with research agents and research papers as two types of nodes, and we consider three types of edges (review, author, and cite) in the graph. Different community activities, such as paper writing and review writing, can be modeled as special message-passing processes on the community graph.
@@
Novel evaluation for research simulation. With ResearchTown for research simulation, a further research question is to evaluate the quality of that. Prior works primarily use human evaluation with breakdown metrics such as novelty, excitement, feasibility, and expected effectiveness (Si et al., 2024; Hu et al., 2024b). These approaches inevitably suffer from subjectiveness and high costs. In our work, since ResearchTown functions as a simulator, our primary focus is on measuring how closely its outputs align with those of the real-world research community. Community graphs naturally provide a similarity-based evaluation method by masking a given paper node in the community graph and evaluating whether a simulator can reconstruct the masked nodes. This definition focuses on simulation similarity, making it scalable and objective. Based on such a node masking prediction task, we build a benchmark called ResearchBench with 1,000 paper writing tasks and 200 review writing tasks requiring multi-agent collaboration.
@@
Main discoveries. Based on the evaluation results from ResearchBench, we highlight three key findings: (1) ResearchTown effectively simulates collaborative research activities, achieving an average similarity score of 0.68 for paper writing and 0.49 for review writing, as measured by the state-of-the-art text embedding model; (2) ResearchTown demonstrates robustness and effectiveness in research simulation, showing improvement when more agents are added and maintaining performance when including unrelated papers; (3) ResearchTown inspires interdisciplinary research, generating innovative ideas that combine insights from NLP, criminology, and astronomy and does not exist in the real-world research.
@@
Stressing ethical concerns. As our work targets simulating the human research community, multiple ethical concerns, including facilitating research plagiarism and producing low-quality or misleading claims, appear. These ethical concerns are addressed in detail in Appendix §B.
@@ 2 Additional Related Work
Graphs with text attributes. In real-world graph tasks, nodes often have textual attributes to carry richer information, forming text-attributed graphs (TAGs) (Yang et al., 2021; He et al., 2023). Previous work on TAGs mainly treats LLMs as tools for understanding text attributes and aims at achieving co-training LLMs and GNNs (Zhao et al., 2023; Chen et al., 2024). In contrast, our approach incorporates agent nodes into the graph, enabling text-based message passing between agent nodes and data nodes. Furthermore, while previous TAG research mainly focuses on node prediction and link prediction tasks (Yan et al., 2023), ResearchTown extends it to both the reconstruction of existing nodes and the prediction of new, non-existent nodes.
@@ 3 Agent-Data Graph for Multi-agent LLMs
Definition of agent-data graphs. To initiate our discussion, we formally define the proposed agent-data graph. An agent-data graph is a special type of heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\mathcal{V}_{a}\cup\mathcal{V}_{d}$ is the node set consisting of two types of nodes, agent nodes and data nodes, and $\mathcal{E}=\mathcal{E}_{aa}\cup\mathcal{E}_{ad}\cup\mathcal{E}_{dd}$ is the edge set consisting of three types of relations, agent-agent, data-data, and agent-data interactions. Here, each data node $v\in\mathcal{V}_{d}$ comes with attributes, e.g., a piece of text, $\mathbf{x}_{v}$; each agent node $u$ is accompanied with an agent function, e.g., an LLM $f_{u}(\cdot)$ with its prompt template and the profile. Each agent function is responsible for two types of tasks: message generation and message aggregation. More details about agent functions are in Appendix §D.3. Without loss of generality, we assume that the data nodes have text attributes, and leave the multi-modal extension of our work, e.g., images, audio, and videos, to future works.
@@
Example of agent-data graphs. Figure 1 shows an example of the agent-data graph. Its definition could be extended to more node types (e.g., codebase, blogs) and edge types (e.g., attend, post, commit). Typically, one blog post can be directly connected to multiple researchers, papers, and other blog posts if they are related to each other.
@@ 4 Building TextGNN on Agent-Data Graphs
$\mathbf{m}_{u}^{(k)}=\textsc{MSG}^{(k)}(\mathbf{h}_{u}^{(k-1)})$ (1)
$\mathbf{h}_{v}^{(k)}=\textsc{AGG}^{(k)}\big{(}\mathbf{h}_{v}^{(k-1)},\{\mathbf{m}_{u}^{(k)}\mid u\in\mathcal{N}(v)\}\big{)}$ (2)
@@
Figure 2: ResearchTown simulation as TextGNN inference on the community graph. The simulation proceeds in three stages: (1) paper reading, where new agent nodes are added based on existing data; (2) paper writing, where data nodes are created; (3) review writing, where the community evaluates and selectively removes (or retains) generated nodes.
Message passing in TextGNN. Following the message-passing process in the standard GNN, we now define a general form of the aggregation function to describe the text-based message-passing process on an agent-data graph $\mathcal{G}$. The key difference between a standard GNN and a TextGNN is that all hidden states in the standard GNN are defined in the embedding space ($\mathbf{h}_{v}\in\mathbb{R}^{d}$) while those in TextGNN are defined in the text space ($\mathbf{h}_{v}\in\Sigma^{*}$). In a TextGNN, we first set the initial hidden states for data nodes $\mathbf{h}_{v}^{(0)}=\mathbf{x}_{v}$, where $\mathbf{x}_{v}$ are text attributes, and the initial hidden states for agent nodes is empty $\mathbf{h}_{u}^{(0)}=\emptyset$. Next, we design a general form of message passing function that handles three distinctive types of interaction, including agent-agent $\mathcal{E}_{aa}$, agent-data $\mathcal{E}_{ad}$, and data-data $\mathcal{E}_{dd}$.
Specifically, the $k$-th TextGNN layer for an agent node $u\in\mathcal{V}_{a}$ can be written as:
$\mathbf{h}_{u}^{(k)}$ $=\textsc{AGG}^{(k)}\big{(}f_{u}(\cdot),\mathbf{h}_{u}^{(k-1)},\{\mathbf{h}_{d}^{(k-1)}\mid(u,d)\in\mathcal{E}_{ad}\},$ (4)
$\quad\{f_{a}(\cdot),\mathbf{h}_{a}^{(k-1)}\mid(u,a)\in\mathcal{E}_{aa}\}\big{)}$
$=f_{u}\Big{(}\Big{[}\mathbf{h}_{u}^{(k-1)},\big{\{}f_{a}\big{(}\big{[}\mathbf{h}_{a}^{(k-1)},\mathbf{h}_{u}^{(k-1)},\mathbf{h}_{d}^{(k-1)}\big{]}\big{)}\mid$
$\quad(u,a)\in\mathcal{E}_{aa},(u,d)\in\mathcal{E}_{ad}\big{\}}\Big{]}\Big{)}$
where []delimited-[][$\cdot$][ ⋅ ] is the concatenation function between texts to fill in the prompt template, $\mathbf{h}_{v}^{(k)}$ represents the hidden states of the $k$-th layer of $v\in\mathcal{V}$, $f_{a}(\cdot)$ represents the agent function paired with the agent node in the neighborhood and $f_{u}(\cdot)$ represents the agent function paired with the agent node.
Similarly, the forwarding process of the $k$-th TextGNN layer for a data node $v\in\mathcal{V}_{d}$ can be written as:
$\mathbf{h}_{v}^{(k)}$ $=\textsc{AGG}^{(k)}\Big{(}\mathbf{h}_{v}^{(k-1)},\{\mathbf{h}_{d}^{(k-1)}\mid(v,d)\in\mathcal{E}_{dd}\},$ (5)
$\quad\{f_{a}(\cdot),\mathbf{h}_{a}^{(k-1)}\mid(v,a)\in\mathcal{E}_{ad}\}\Big{)}$
$=f_{g}\Big{(}\Big{[}\mathbf{h}_{v}^{(k-1)},\big{\{}f_{a}\big{(}\big{[}\mathbf{h}_{a}^{(k-1)},\mathbf{h}_{v}^{(k-1)},\mathbf{h}_{d}^{(k-1)}\big{]}\big{)}\mid$
$\quad(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\big{\}}\Big{]}\Big{)}$
@@ 5 ResearchTown: Applying TextGNN to Community Graph
Inputs and outputs of ResearchTown. Building on the definitions of TextGNN and the agent-data graph in Section §3 and Section §4, we simulate different research activities by modeling each as a specific instantiation of a TextGNN layer. ResearchTown processes diverse research materials and produces structured outputs. The input varies by task: only paper abstracts are used for paper reading and writing, while full papers are provided for review writing. The output format is also task-specific: paper reading generates profile descriptions, paper writing generates bullet-point summaries, and review writing produces bullet-point critiques along with a numerical review score. These standardized output formats—described in more detail in Appendix §F—facilitate evaluation over long-context inputs and enable fine-grained, sub-component similarity scoring.
Hidden states of ResearchTown. In ResearchTown, the hidden state of each node represents a condensed version of research materials, such as papers or reviews. Initially, paper nodes are initialized with the full text of papers. Through iterative message passing, these nodes gradually evolve into a standardized bullet-point format, distilling key information for easier downstream evaluation. Similarly, review attributes associated with paper nodes are also represented using bullet points to make it in a compact form. Bullet-point compact form with limited length allows TextGNN to conduct message passing multiple times efficiently.
@@
Agent-data graph for research community modeling - community graph. We adopt the agent-data graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to research community simulation, which we named as community graph. As is shown in Figure 2, each agent node $\mathcal{V}_{a}$ represents one researcher, and each data node $\mathcal{V}_{d}$ represents a paper. The edge set $\mathcal{E}_{dd}$ captures paper citations, the edge set $\mathcal{E}_{ad}$ captures authorship (a researcher writes a paper) and reviewing expertise (a researcher is qualified to review a paper). We omit the edge set $\mathcal{E}_{aa}$ to simplify the framework, as a collaboration between authors can typically be inferred through 2-hop paths via $\mathcal{E}_{ad}$ edges.
@@
TextGNN for research activity simulation. Based on the constructed community graph, we further identify the key types of research activities where TextGNN can be used for simulation. Specifically, as shown in Figure 2, we split the research simulation process into three critical stages: (1) paper reading, (2) paper writing, and (3) review writing. We believe these stages are crucial in the research community, and each stage relies on the output of the previous stage as input. We provide a detailed description for each stage and the corresponding TextGNN layer definition below.
@@
$\triangleright$ Stage 1: Paper reading. Reading papers to collect insights is a necessary process for initializing a research project. In the community graph, the paper reading process can be described as inserting a new agent node to the community graph and aggregating its neighborhood information based on Equation 4. Here, the new agent profile is non-existent before reading a collection of papers, and the profile is created after the paper reading process, making the TextGNN layer unique. Concretely, by adapting Equation 4, the TextGNN layer for paper reading can be written as:
@@
$\begin{split}\mathbf{h}_{u}&=\textsc{AGG}\Big{(}f_{u}(\cdot),\{\mathbf{h}_{d}\mid(u,d)\in\mathcal{E}_{ad}\}\Big{)}\\ &=f_{u}\Big{(}\Big{[}\left\{\mathbf{h}_{d}\mid(u,d)\in\mathcal{E}_{ad}\right\}\Big{]}\Big{)}\end{split}$ (6)